# Stability of fully Discrete Local Discontinuous Galerkin method for the generalized Benjamin-Ono equation

### v5

@@ 3.1. Stability analysis of Crank-Nicolson fully discrete LDG scheme
To develop the fully discrete local discontinuous Galerkin (LDG) scheme, we utilize the Crank-Nicolson method for time discretization in (2.8). We partition the time domain using a time step $\tau$. Let $\{t_{n}=n\tau\}_{n=0}^{M}$ be the partition of the given time interval $[0,T]$, and define $u^{n}=u(t_{n})$ and $u^{n+\frac{1}{2}}=\frac{u^{n+1}+u^{n}}{2}$.
@@ 3.2. Stability analysis of higher order fully discrete LDG scheme
Hereby we focus on the stability analysis with an explicit fourth-order Runge-Kutta time discretization. The spatial stability of LDG schemes are demonstrated in Section 2. It is worth mentioning that the stability analysis under the higher-order time discretizations is a challenging task, and at present we have pursued this analysis exclusively in the linear case. In particular, for simplicity, we choose the zero flux function for the subsequent analysis [24]. We would like to remark that to the best of our knowledge, the stability analysis for fourth-order LDG schemes involving general LDG operators has not been performed in the literature. The following analysis can be considered as a contribution to the stability analysis under the aforementioned assumptions.
@@
Since the LDG operator is semi-negative, applying the two-step stability result from [21, Theorem 2.1], we have
@@ 4. Numerical Experiments
In this section, our goal is to validate the proposed LDG scheme (2.7) for the Benjamin-Ono equation (1.1). We begin our numerical validation using the Crank-Nicolson (CN) LDG scheme (3.1) with $\tau=0.5h$ and $k=1$. We employ the periodic one soliton solution of the Benjamin-Ono equation. It is important to note that since the implicit term appears in the nonlinear part of the scheme (3.1), we use the Newton iteration to solve it at each time step.
@@
We compare approximate solution $u$ obtained by (3.1) with the exact periodic solution of the Benjamin-Ono equation (1.1) with $f(U)=\frac{1}{2}U^{2}$. For instance, we consider the solution of (1.1) from Thomée et. al. [23] (also see [9, 10])
@@
Our focus lies on verifying the performance of the scheme using a low storage explicit Runge-Kutta (LSERK) of fourth-order time discretization [4, 11, 13] of the form
@@
where the weighted coefficients $a_{j}$, $b_{j}$ and $c_{j}$ of the LSERK method given in [13]. The above defined iteration is equivalent to the classical fourth-order method (3.4) [13]. LSERK is considerably more efficient and accurate than (3.4) since it has the disadvantage that it requires four extra storage arrays. We are providing the steps of algorithm to compute the approximate solution using LDG scheme (2.7) with LSERK time discretization, for more details one may refer to [13].
Algorithm 1 Numerical Procedure for Benjamin-Ono Equation
1:  Set up problem parameters and initial conditions
2:  Compute time step size $dt$ based on CFL condition
3:  Determine total number of time steps $Nsteps$ based on Final time and $dt$
4:  for $tstep=1$ to $Nsteps$ do
5:     for $INTRK=1$ to $5$ do
6:        Compute local time $timelocal$ for current stage
7:        Evaluate right-hand side $Lu$ of the differential equation using Algorithm 2
8:        Update solution $u$ using Runge-Kutta stages
9:     end for
10:     Increment time by $dt$
11:  end for
12:  return  Final solution $u$
Algorithm 2 Algorithm for Computing RHS of equation (4.2)
0:  Solution $u$, left boundary value $xL$, right boundary value $xR$, current time $timelocal$
0:  Right-hand side of the equation $rhsu$
1:  Compute coefficients and variables based on initial conditions and boundary values
2:  Define field differences at interfaces, including boundary conditions
3:  Compute flux for $u$
4:  Compute local variable $q$ and its differences using $u$
5:  Compute local variable $p$, using Hilbert transform of $q$
6:  Compute flux for $p$ ($fluxp$)
7:  Compute squared differences of $u$, $du^{2}$ at interfaces, and impose boundary conditions
8:  Compute flux for the convective term ($fluxf$)
9:  Compute derivative of the flux ($dfdr$)
10:  Compute right-hand side of the semi-discrete PDE ($rhsu$)
11:  return  $rhsu$
@@
We have also approached to compute the approximation of the Hilbert transform by using the Gaussian quadrature for other soliton solutions [12, 23]. However, we do not achieve the optimal rates or the theoretical rates obtained by convergence analysis. Similar observations are also described in [12, Section 5]. One possible interpretation is that the principal value integral is not being computed accurately near the singularity, and it is most likely caused by our approximation of the full line problem by restricting to a bounded interval while still employing the full line Hilbert transform.
@@ 5. Concluding remarks
We have designed a local discontinuous Galerkin method for the Benjamin-Ono equation with general nonlinear flux. We have shown that the semi-discrete scheme is stable. The stability analysis is also carried out for the fully discrete Crank-Nicolson LDG scheme considering any general nonlinear flux function and strong stability of fourth-order Runge-Kutta LDG scheme under certain assumptions on the flux function. The theoretical convergence analysis for fully discrete scheme incorporating fourth-order Runge-Kutta time marching scheme involving general nonlinear flux will be addressed in the future work. In the numerical experiments, it is observed that the $L^{2}$-error is quite small even for the coarser grids and the optimal rates are obtained for various degrees of polynomials which demonstrates the efficiency and accuracy of the proposed scheme. In addition, the proposed fully discrete schemes preserve the conservative quantities such as mass and momentum.
@@
The authors declare that they have no known competing financial interests that could have appeared to influence the work reported in this paper. Furthermore, no data was used for the research described in the article.

### v6

@@ 3.1. Stability analysis of Crank-Nicolson fully discrete LDG scheme
To develop the fully discrete local discontinuous Galerkin (LDG) scheme, we utilize the Crank-Nicolson method for time discretization in (2.8). We partition the time domain using a time step $\tau$. Let $\{t_{n}=n\tau\}_{n=0}^{M}$ be the partition of the given time interval $[0,T]$, and define $u^{n}=u(t_{n})$ and $u^{n+\frac{1}{2}}=\frac{1}{2}(u^{n+1}+u^{n})$.
@@ 3.2. Stability analysis of higher order fully discrete LDG scheme
Hereby we focus on the stability analysis with an explicit fourth-order Runge-Kutta time discretization. The spatial stability of the LDG scheme is demonstrated in Section 2. It is worth mentioning that the stability analysis under the higher-order time discretizations is a challenging task, and at present we have pursued this analysis exclusively in the linear case. In particular, for simplicity, we choose the zero flux function for the subsequent analysis [26]. We would like to remark that to the best of our knowledge, the stability analysis for fourth-order LDG schemes involving general LDG operators has not been performed in the literature. The following analysis can be considered as a contribution to the stability analysis under the aforementioned assumptions.
@@
Since the LDG operator is semi-negative, applying the two-step stability result from [23, Theorem 2.1], we have
@@ 4. Numerical Experiments
In this section, our goal is to validate the proposed LDG scheme (2.7) for the Benjamin-Ono equation (1.1). We begin our numerical validation using the Crank-Nicolson (CN) LDG scheme (3.1) with $\tau=0.5h$ and $k=1$. We employ the periodic one soliton and two soliton solutions of the Benjamin-Ono equation. It is important to note that since the implicit term appears in the nonlinear part of the scheme (3.1), we use the Newton iteration to solve it at each time step.
@@
Example 1
We compare the approximate solution $u$ obtained by (3.1) with the exact periodic solution of the Benjamin-Ono equation (1.1) with $f(U)=\frac{1}{2}U^{2}$. For instance, we consider the solution of (1.1) from Thomée et al. [25] (also see [10, 11])
@@
Our focus lies on verifying the performance of the scheme using a low storage explicit Runge-Kutta (LSERK4) of fourth-order time discretization [4, 12, 14] of the form
@@
where the weighted coefficients $a_{j}$, $b_{j}$ and $c_{j}$ of the LSERK4 method are given in [14]. The above defined iteration is equivalent to the classical fourth-order method (3.4) [14]. LSERK is considerably more efficient and accurate than (3.4) since it has the disadvantage that it requires four extra storage arrays. To compute the approximate solution using the LDG scheme (2.7) with LSERK time discretization, one may refer to [14].
@@
Example 2
We consider a two-soliton solution $U_{2}(x,t)$ to the Benjamin-Ono equation (1.1), derived using the inverse scattering transform method. The expression is given by [25, 13]:
$U_{2}(x,t)=\frac{4c_{1}c_{2}\left(c_{1}\lambda_{1}^{2}+c_{2}\lambda_{2}^{2}+\frac{(c_{1}+c_{2})^{3}}{c_{1}c_{2}(c_{1}-c_{2})^{2}}\right)}{\left(c_{1}c_{2}\lambda_{1}\lambda_{2}-\frac{(c_{1}+c_{2})^{2}}{(c_{1}-c_{2})^{2}}\right)^{2}+(c_{1}\lambda_{1}+c_{2}\lambda_{2})^{2}},$ (4.3)
where $\lambda_{1}$ and $\lambda_{2}$ are given by
$\lambda_{1}=x-c_{1}t-d_{1},\quad\lambda_{2}=x-c_{2}t-d_{2},$
and we choose the parameters $c_{1}=0.3,\,c_{2}=0.6,\,d_{1}=-30,\,d_{2}=-55.$ The LDG formulation involves evaluation of terms of the form $(\mathcal{H}\phi_{i},\psi_{i})$ for basis functions $\phi_{i},\psi_{i}$ in each element $I_{i}$. These integrals are computed using the Gaussian quadrature. For the first Hilbert integral, we use a 7-point Gaussian quadrature in each element $I_{i}$. For the second Hilbert integral involving interaction terms, we use an 8-point Gaussian quadrature. This ensures high accuracy in evaluating the nonlocal dispersive contribution while retaining efficiency.
We set the initial data as $u_{0}(x)=U_{2}(x,0)$ and run the simulation until final time $T=180$. We compare the numerical solution $u(x,T)$ with the exact two-soliton profile $U_{2}(x,T)$ using $L^{2}$-norm error metrics. The mesh consists of $N$ uniform elements and we test polynomial degrees $k=1,2,3$. We employ the LSERK4 scheme to achieve high-order accuracy in time while maintaining computational efficiency.
Table 4.4 confirms that the proposed LDG scheme achieves the optimal convergence rates of order $k+1$ in the $L^{2}$-norm for polynomial degree $k=1,2,3$, verifying the accuracy of the spatial discretization. Figure 4.2 shows the comparison of the exact and numerical two-soliton convergence at $T=180$ for $k=3$ with $N=2560$. The LDG scheme successfully captures the nonlinear interaction and the propagation of both solitons with high accuracy.
The numerical experiment confirms the accuracy and efficiency of the LDG scheme for solving the generalized Benjamin-Ono equation. By carefully selecting two-soliton parameters, we validated that the scheme accurately captures nonlinear wave interactions. Both Crank–Nicolson and LSERK4 time discretizations yield optimal rates of convergence. The high-order Gaussian quadrature used for evaluating the Hilbert transform further ensures the fidelity of the method.
$k$ $k=1$ $k=2$ $k=3$
$N$ $E$ $R_{E}$ $E$ $R_{E}$ $E$ $R_{E}$ $\mathbf{C_{1}^{h}}$ $\mathbf{C_{2}^{h}}$
640 5.82e-01 4.92e-02 3.86e-03 0.997 0.984
2.08 2.77 3.93
1280 1.36e-01 7.36e-03 2.47e-04 0.999 0.998
2.01 2.91 3.97
2560 3.38e-02 9.71e-04 1.57e-05 1.000 1.000
2.00 2.95 3.92
5120 8.44e-03 1.23e-04 1.03e-06 1.000 1.000
Table 4.4. $L^{2}$-error and convergence rate $R_{E}$ for the LDG scheme with LSERK4 time integration at $T=180$ for the two-soliton solution, using $N$ elements and polynomial degree $k$. Normalized conserved quantities $\mathbf{C_{1}^{h}}$ and $\mathbf{C_{2}^{h}}$ confirm numerical mass and energy conservation.
Figure 4.2. Exact $U_{2}(x,t)$ and approximate solution $u(x,t)$ computed by the RK-LDG scheme at $T=180$ with $N=2560$, $k=3$ and initial condition $U_{2}(x,0)$ of (1.1).
@@ 5. Concluding remarks
We have designed a local discontinuous Galerkin method for the Benjamin-Ono equation with general nonlinear flux. We have shown that the semi-discrete scheme is stable and obtained a suboptimal order of convergence. The stability analysis is also carried out for the fully discrete Crank-Nicolson LDG scheme considering any general nonlinear flux function and strong stability of the fourth-order Runge-Kutta LDG scheme under certain assumptions on the flux function. The theoretical convergence analysis for the fully discrete scheme incorporating the fourth-order Runge-Kutta time marching scheme involving general nonlinear flux will be addressed in future work. In the numerical experiments, it is observed that the $L^{2}$-error is quite small even for the coarser grids and the optimal rates are obtained for various degrees of polynomials which demonstrate the efficiency and accuracy of the proposed scheme. In addition, the proposed fully discrete schemes preserve the conservative quantities such as mass and momentum.
@@
The authors declare that they have no known competing financial interests that could have appeared to influence the work reported in this paper. Furthermore, no data were used for the research described in the article.